### v1

# Divide et Impera: Learning impedance families for peg-in-hole assembly

Abstract
This paper addresses robotic peg-in-hole assembly using the framework of Elementary Dynamic Actions (EDA). Inspired by motor primitives in neuromotor control research, the method leverages three primitives: submovements, oscillations, and mechanical impedances (e.g., stiffness and damping), combined via a Norton equivalent network model. By focusing on impedance parameterization, we explore the adaptability of EDA in contact-rich tasks. Experimental results, conducted on a real robot setup with four different peg types, demonstrated a range of successful impedance parameters, challenging conventional methods that seek optimal parameters. We analyze our data in a lower-dimensional solution space. Clustering analysis shows the possibility to identify different individual strategies for each single peg, as well as common strategies across all pegs. A neural network model, trained on the experimental data, accurately predicted successful impedance parameters across all pegs. The practical utility of this work is enhanced by a success-predictor model and the public availability of all code and CAD files. These findings highlight the flexibility and robustness of EDA; show multiple equally-successful strategies for contact-rich manipulation; and offer valuable insights and tools for robotic assembly programming.
Index Terms:
Robotic assembly, Physical interaction, Impedance control, Neural networks.
I Introduction
A current trend in robotic control is the development of contact-rich manipulation techniques [1, 2, 3, 4, 5, 6]. Despite the existence of many effective methods for robot motion planning, managing the physical interaction of robots with various objects and surfaces remains a significant research challenge. However, mastering contact-rich scenarios is crucial for robots to achieve human-like interaction capabilities, such as those required in industrial assembly lines.
Lately, robotics has seen promising developments in contact-rich manipulation, such as model-based methods [7, 8, 5] and learning-based approaches [9, 3, 10, 6]. Model-based methods leverage mathematical representations of contact dynamics, ensuring that the robots can predict and react to contact forces accurately. Learning-based approaches utilize probabilistic models to handle uncertainties in contact interactions. Learning approaches, particularly those based on reinforcement learning and neural networks, enable robots to learn optimal contact strategies through trial and error, leading to adaptable manipulation skills.
Contact-model-based optimization approaches integrate detailed contact models into the control framework, considering factors such as friction cones, contact modes, and complementarity constraints. These models are crucial for tasks requiring precise interaction with objects, such as manipulation in cluttered environments. Techniques like Mixed Integer Quadratic Programming, which treat contact modes as integer variables [11, 12], and Linear Complementarity Problems [13, 7, 14], are commonly used. However, these methods face significant challenges, including numerical complexity, sensitivity to coordinate frame choice and object geometry, and the need for accurately estimated parameters like friction coefficients.
Sampling-based optimization methods have also gained traction, particularly in scenarios where gradient information is not readily available. Most of these sampling-based methods are Zeroth-order methods, which do not rely on gradients and sample multiple actions and update control parameters based on outcomes without needing derivative information [15]. Variations such as STOMP [16], PI² [17, 18], and MPPI [19, 20, 21] have become popular. First-order methods, involving randomized smoothing for gradient-based optimization, are also employed [22, 5]. However, these methods are often constrained by the choice of the cost function, which will only output one set of control parameters, even though multiple solutions might be feasible. Moreover, those methods often lead to trivial solutions that merely resemble position controllers [17, 23].
Other common approaches in robotic control are Dynamic Movement Primitives (DMP) and Elementary Dynamic Actions (EDA). Both approaches are inspired by human motor control research [24, 25, 26, 27] and try to leverage insights into how humans achieve complex, coordinated movements [28, 29]. It is hypothesized that keys to human performance are building blocks of coordination, as originally proposed by Sherrington [30]. These building blocks are conventionally called “motor primitives” and are basic motor actions that can generate a rich repertoire of behaviors. Bernstein further expanded on this concept by introducing the idea of “repetition without repetition” [31]. He observed that humans can accomplish tasks in numerous ways, adapting to subtle environmental changes and learning strategies for movement. This adaptability is a key ability of the human motor system. Latash argued that the motor system does not operate with explicit costs but rather utilizes a subspace of multiple solutions, enabling flexible and effective movement strategies [32]. Wolpert et al. emphasized the importance of motor learning, highlighting that feed-forward control is essential for natural movements, as feedback mechanisms can be too slow[33, 34].
The main motivation of this work is to investigate whether we can achieve adaptability comparable to that observed in human motor control studies by using a “human motor control inspired” robot controller.
In this paper, our main method will be EDA, originally introduced by [26, 27]. The key to this approach is breaking down complex physical interaction tasks into simpler, fundamental actions that can be more easily controlled and adapted. EDA consists of three distinct classes of motor primitives: submovements and oscillations as kinematic primitives [35], and mechanical impedances as distinct primitives to manage physical interaction [36, 26, 37]. The three primitives can be combined using a Norton equivalent network model [38]. More details about the primitives of EDA and the Norton equivalent network model can be found in appendix A.
While EDA has demonstrated remarkable success in contact-rich manipulation [39, 27, 40], parameterizing the distinct primitives remains a key challenge. This paper aims to elucidate the choices of parameters for interaction primitives, specifically impedance (e.g., stiffness and damping). Our use-case will focus on peg-in-hole assembly which is a common task in robotic research and industrial assembly (see fig. 1). To our knowledge, this paper is the first to go beyond generating a controller for contact-rich manipulation, but aims to illuminate the structure, effects, and distribution of successful parameter choices for peg-in-hole assembly. Our goal is to provide novel insights that can enhance our understanding and application of EDA in practical scenarios, as well as to help in finding more intuitive solutions to complex manipulation tasks.
Figure 1: Exemplary peg-in-hole assembly using a KUKA LBR robot. All control parameters were defined with respect to the spatially fixed coordinate frame, highlighted in orange.
The main contributions of this paper are to investigate four distinct claims. The knowledge gained from this investigation will culminate in a success predictor for robotic peg-in-hole assembly, which is intended to help less-experienced programmers to use our method.
  • Claim 1: We demonstrate that Elementary Dynamic Actions (EDA) do not restrict successful assembly to a single optimal set of parameters but instead enable a range of successful impedance parameters. This finding challenges the conventional method of finding a single optimal solution in robotic assembly tasks.
  • Claim 2: Our analysis reveals the existence of a non-arbitrary, lower-dimensional representation of successful assembly strategies. This simplification is achieved through Principal Component Analysis (PCA), which shrinks the parameter space dimension from eight to three without substantially removing information.
  • Claim 3: We identify that different assembly strategies are effective for individual peg types. This highlights the adaptability of EDA to accommodate variations and provide a tailored approach to different assembly challenges.
  • Claim 4: Across various pegs, common assembly strategies emerge, suggesting that impedance settings may be generalized across different tasks. This would support the development of more streamlined and efficient programming practices.
  • Application: We introduce a neural network model as a success-predictor for peg-in-hole assembly. This model offers a practical tool to predict assembly outcomes based on impedance parameters and can serve as a valuable resource for less-experienced robot programmers.
All hardware and software tools used in this work are publicly available, to enable researchers to test our findings, extend our dataset, or even use the software for real peg-in-hole assembly applications in industry.
II Methods
This section outlines the theoretical framework and experimental protocol used in this work to investigate the role of impedance parameter choices in robotic peg-in-hole assembly. We employed a systematic approach to evaluate the role of EDA and their components to achieve robotic assembly operations. All methods introduced here were used to investigate the main claims of the paper.
II-A Assembly strategy for peg-in-hole tasks
For our assembly application, we used a KUKA LBR robot and the Cartesian Impedance Controller of the KUKA Sunrise framework 1.17. Details of the algorithms can be found in [41]. The robot application code is provided on Github [^1] https://github.com/mosesnah-shared/ImpedanceClassification/tree/main/Sunrise.
We programmed an application to assemble four different peg-in-hole assemblies: square, triangular, hexagonal, and cylindrical pegs. Between the workpieces and the respective insertion holes was a nominal clearance of 0.14 mm (square peg), 0.18 mm (hexagonal peg), and 0.20 mm (circular and triangular pegs). All pieces were printed with a PRUSA i3 MK3 3D-printer using PLA filament. All components except the pegs were printed with $20\%$ infill with gyroid pattern. Pegs were printed with higher infill ($>70\%$) to guarantee structural strength. The CAD data of all pieces can be found on Dropbox [^2] https://www.dropbox.com/scl/fo/ljrkha4bh17v4scy8uy9l/AL4zqjOkfwap6kvq8QBylWQ?rlkey=qqyxf8euipigzhlokfo5x2nr6&dl=0.
Algorithm 1 shows the structure of the assembly application. For our application, we used the torque sensors of the robot to measure external contact (Algorithm 1, line 3-5 and 14-20).
Algorithm 1 Primitives for peg-in-hole assembly: submovements (orange), oscillations (blue), and mechanical impedance (green). The coordinates related to the algorithm can be seen in fig. 1. The impedance parameters in line 11-13 are exemplary and were varied throughout the experiments. The frame-based discrete movements and oscillations remained fixed throughout the experiments.
1:Incline workpiece and align planarly at center of hole
2:Create spatial Force Condition: $F_{\text{TCP}}<15$ N
3:while Force Condition do
4:     Move linRel along $-x$ and $y$
5:end while
6:Rotate from current TCP-pose to final TCP-pose, with
7:      Siffness:
8:      CartDOF.TRANSL $=400$ [N/m]
9:     //Be compliant about A
10:   //for peg to ‘‘slip’’ into the hole
11:      CartDOF.A $=5$ [Nm/rad]
12:      CartDOF.B,C $=100$ [Nm/rad]
13:      Damping CartDOF.ALL $=0.7$
14:while  Force Condition  do
15:     Move linRel along $-z$, with
16:     Siffness (line 8-12), Damping (line 13)
17:      //Handle riples of 3D-print
18:     Overlay Oscillation about CartDOF.A, with
19:     Amplitude $A=0.6$ [N], Frequency $f=4.5$ [Hz]
20:end while
The overall assembly strategy was inspired by observations of humans performing assembly tasks, often without visual feedback. Humans frequently rely on tactile feedback and subtle movements, such as wiggling the workpiece, to align and insert components accurately. All primitives of EDA were used. For the kinematic primitives (submovements and oscillations), the frame-based programming procedure of KUKA Sunrise was used. For the impedance parameters, the diagonal terms of the translational and rotational stiffness matrices and the translational and rotational damping ratios were varied throughout the experiments.
With the assembly strategy (Algorithm 1), the peg automatically “slipped” into the hole, even though the workpiece was not perfectly aligned during the assembly [42]. This is especially beneficial for end-of-line assembly where the workpiece is not perfectly positioned in front of the robot. Expensive fixation mechanisms can be obviated.
Due to imperfections of the 3D-print, the rippled surfaces of the hole and the workpiece could impede the insertion of the peg. We therefore overlaid small oscillations on the discrete motion along $-z$ during the assembly (Algorithm 1, line $18-19$). While certainly other assembly strategies would be successful for this application (e.g., combining position control and perception), for us the key was the combination of tactile exploration with kinematic primitives and impedance. This demonstrated the benefits of the modularity of EDA. A video of the assembly strategy can be found on YouTube [^3] https://youtu.be/T2pFu6IXk04.
As can be seen in algorithm 1, many parameters must be selected. The exact number of parameters depends on the assembly strategy of the programmer. In this paper, we emphasized the impedance parameter choice (green). Thus, for our experiments we fixed the parameters of the kinematic primitives (submovements and oscillations in algorithm 1). With this, a set of eight impedance values remained to be selected. The scalar stiffnesses along and about the coordinates $\{x,y,z\}$ are denoted as $\{k_{x},k_{y},k_{z}\}$ and $\{k_{C},k_{B},k_{A}\}$, respectively. Note that the notation is defined in this way by the KUKA Sunrise programming interface. For example, $k_{C}$ represents the rotation about the $x$-axis . The common damping ratio values for translational and rotational motion are denoted $\{\zeta_{t},\zeta_{r}\}$. More information about the controller parameterization can be found in the KUKA Sunrise manual [^4] https://my.kuka.com/s/product/kuka-sunriseos-117/.
II-B Experimental protocol
Figure 2: Experimental setup used during the trials. The workpiece with the hole was secured to prevent any movement throughout the four 40-hour long experimental trials.
The control parameters that can realistically be rendered by the robot were determined to be in a range of $50\leq k_{t}\leq 1000$ N/m for translational stiffnesses, $5\leq k_{r}\leq 200$ Nm/rad for the rotational stiffnesses, and $0.1\leq\zeta\leq 0.9$ for damping ratio. These ranges were found empirically.
Finding feasible impedance values is not trivial. Our intention was to test combinations of impedance values for peg-in-hole assembly. Due to the curse of dimensionality [43], it was not practical to test all possible combinations of impedance values. Hence, we assigned three different categories for each of the parameter ranges: small, medium, and large. The exact values of the three categories can be found in appendix B.
For each impedance parameter, we randomly selected a value from one of the three categories totalling $65613^{8}=$ 6561 impedance combinations tested for each peg (fig. 2). As each iteration required an average of 22 seconds, the total experimental time for each peg took about 40 hours. For four pegs, this gave an overall experimental time of about 160 hours.
After the experiments, we investigated four different claims:
  • Claim 1: Success with a range of parameters
  • Claim 2: Lower dimensional representation of solutions
  • Claim 3: Different assembly strategies for individual pegs
  • Claim 4: Common assembly strategies across all pegs
The methods used to evaluate these claims are explained in detail in the following.
II-C Claim 1: Success with a range of parameters
II-C1 Success Rate
A total of $N_{tot}=3^{8}=6561$ impedance parameter combinations was tested for each peg shape. The success rate, $s_{r}$, was computed as the percentage of successful trials, $N_{s}$, over the number of total tested parameters, $N_{tot}$.
$s_{r}=100\cdot\frac{N_{s}}{N_{tot}}$ (1)
II-C2 Linear regression model
To investigate Claim 1—that EDA enables a range of successful parameters, rather than one optimal set—we used linear regression models to fit the success rate (dependent variable) as a function of the impedance parameters (independent variables) for each peg. The goodness-of-fit for each model was assessed using the coefficient of determination ($R^{2}$), which measures the variance in the dependent variable that is predictable from the independent variable. The statistical significance of the regression slope was tested against the null hypothesis of no relation between the success rate and the impedance parameters examined. The results were interpreted to understand how different impedance parameter choices affected the outcome of the peg-in-hole task. For example, a significant positive slope in the relationship between success rate and the stiffness $k_{z}$ implied that higher vertical stiffness tended to improve task success.
II-D Claim 2: Lower dimensional representation of solutions
II-D1 Principal Component Analysis
Principal Component Analysis (PCA) is a statistical technique used to reduce the complexity of high-dimensional data by projecting it onto a smaller number of dimensions, known as principal components (PCs). This technique is particularly useful for revealing hidden patterns and explaining variance in data.
For our study, we applied PCA to the impedance solutions obtained from our peg-insertion task to test if a non-arbitrary lower-dimensional subspace could account for our data (Claim 2). We projected the data into a subspace defined by the PCs that captured the most variance. The main PCs provide insight into the principal directions in the data. Moreover, a significant reduction in dimension, from eight to three or fewer, would enable us to visualize the 8-dimensional solution space in the lower-dimensional PC space. Specifically, to extract PCs from the impedance solutions, we utilized the pca function from MATLAB 2022b.
As PCA is data-driven, it may identify a lower degree-of-freedom space in random data. To ensure that any reduction of the solution space observed here was not a random artifact, we compared the slope of the variance-accounted-for by the first three dominant PCs to the slope of the variance-accounted-for by an equally-sized random subset of trials. We conducted a pair-wise t-test between slopes of the variance-accounted-for in the successful trials and the random trials. If the slopes in the successful trials were significantly steeper (i.e., the probability that the slopes were the same was $\alpha\leq 0.05$), we concluded that the PCs extracted from the successful trials were not a random artifact. Specifically, we could claim that there existed a non-arbitrary impedance solution subspace.
Additionally, we projected both successful and unsuccessful trials into the identified PC subspace to visually assess their distribution. If unsuccessful trials consistently fell outside the identified subspace, this would further support the non-arbitrary nature of the impedance solution subspace.
II-E Claim 3: Different assembly strategies for individual pegs
II-E1 K-means Clustering
To investigate Claim 3—that families of solutions encode different assembly strategies—-we utilized k-means clustering on successful trials within our 8-dimensional impedance solution space. The identified clusters would represent distinct common impedance solution strategies. We employed MATLAB 2022b’s k-means clustering algorithm with an online phase update and 5 replicates [44, 45]. The online updates, wherein each sample was individually reassigned to a new cluster if it reduced the sum of point-to-centroid distances, ensured convergence to a local minimum. The use of multiple replicates added robustness to the clustering process.
To identify the effective number of clusters we computed the silhouette value $s_{i}$ of each data point using the equation
$s_{i}=\sum_{i=1}^{N}\frac{b_{i}-a_{i}}{max(a_{i},b_{i})},$ (2)
where $i$ is the $i-$th data point, $a_{i}$ is the average distance between the $i-$th point and all other data points within the cluster, and $b_{i}$ is the avg distance between the $i-$th point and all other data points within the next nearest cluster. From the above definition it is clear that $-1\leq s_{i}\leq 1$. Values close to 1 indicate well-clustered data, negative values suggest misclustered data, and values near zero indicate the datum is near the border of two clusters [46, 47].
The mean $s_{i}$ over all points of a cluster is a measure of how tightly grouped all the points in the cluster are. Thus, the mean $s_{i}$ over the entire dataset is a measure of how appropriately the data have been clustered. If there are too many or too few clusters, as may occur when a poor choice of $k$ is used in the clustering algorithm, some of the clusters will typically display much lower silhouette values than the rest.
With this, the overall clustering effectiveness was evaluated by the silhouette score, $SC$, defined as:
$SC=\max_{k}s(k),$ (3)
where $k$ represents the number of clusters and $s(k)$ represents the mean over all $s_{i}$ of the entire dataset for that number of clusters $k$. The number of clusters $k$ that maximizes the silhouette coefficient $SC$ is the most effective number of clusters for the k-means clustering algorithm [46, 47].
It is important to note that k-means clustering must have at least two clusters and can have at most as many clusters as the number of data points provided. In the latter case, the silhouette coefficient will approach 1 as each point provides its own cluster. Visual representation was facilitated by projecting the impedance solution data for each peg onto the PCA space established previously.
We conducted a one-way MANOVA to determine if the identified clusters were distinct. Detailed methodology can be found in appendix C.
II-F Claim 4: Common assembly strategies across all pegs
II-F1 Neural Networks for Binary Classification
To investigate Claim 4—that common assembly strategies applied to all pegs—we used a neural network for binary classification. Our goal was to find the best-fit model that predicted success for our experimental data with $N=4*3^{8}$ data points and labels $y=\{0,1\}$, i.e., true (1) and false (2). The “best-fit” was determined by the minimum of the Binary Cross Entropy:
$BCE=-\sum^{N}_{i=1}\{y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\}.$ (4)
We used a binary cost, as only success or failure is important for industrial peg-in-hole assembly, not the minimization of a continuous error function such as the Euclidean error.
For given input impedance parameters $\bm{x}=\{k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}\}$, we wanted to find the function $f(\bm{x})$ that represents the boundary of our data, dividing it into success or failure. For the neural network, we chose the following structure with two hidden layers:
$f(\bm{x})=f_{3}(\bm{W}_{3}\ f_{2}(\bm{W}_{2}\ f_{1}(\bm{W}_{1}\bm{x}+\bm{b}_{1})+\bm{b}_{2})+b_{3}).$ (5)
Here, $\bm{W}_{1},\bm{W}_{2},\bm{W}_{3}$ are the weights of the network, $f_{1},f_{2},f_{3}$ are activation functions, and $b_{1},b_{2},b_{3}$ are bias terms associated with each neuron in the network.
The network was trained for all pegs, i.e., the eight impedance values and their success flags were merged into one dataset. Once the network was trained, we used the network to produce successful data and analyze whether all pegs shared a common assembly strategy, represented by a lower-dimensional PC space (sec. III-D).
The two hidden layers and output of the neural network were defined as follows: the first hidden layer transformed the eight-dimensional normalized input features $\mathbf{X}^{(i)}=\{k_{x}^{(i)},k_{y}^{(i)},k_{z}^{(i)},k_{C}^{(i)},k_{B}^{(i)},k_{A}^{(i)},\zeta_{t}^{(i)},\zeta_{r}^{(i)}\}\in\mathbb{R}_{>0}^{8}$ to 64 neurons; the second hidden layer took the 64 features from the hidden layer and passed them to another 64 neurons; the output layer transformed the 64 outputs of the second layer to the predicted output $\hat{y}^{(i)}$. After each of the first two layers, a Rectified Linear Unit (ReLU) activation function was used to introduce non-linearity into the model. The Sigmoid activation function was used at the output layer to map the final value to a probability between 0 and 1. The Adam optimizer [48] was used to update the weights of the network with a fixed learning rate of $1\times 10^{-4}$.
III Results
III-A Claim 1: Success with a range of parameters
III-A1 Success Rate
TABLE I: Percentage success rate and total number of successes for the different tested pegs.
Peg Success Rate [%] Total Success
Triangle 1.54 101/6561
Square 9.85 646/6561
Hexagon 2.76 181/6561
Cylinder 2.24 147/6561
Tab. I reports the success rates for every peg as well as the absolute number of total successes. Excluding the square, which presented the highest success rate ($9.9\%$), the other three pegs presented a success rate ranging between $1.5\%$ and $2.8\%$. We hypothesize that the success rate of the square peg was substantially higher because we used this peg for the initial programming of the application, which determined the kinematic primitives (translations, rotations, and oscillation) and were unchanged for the other pegs (sec. II). In an absolute sense, among the $6561$ tested impedance parameters, it was possible to find between $101$ (triangle) and $646$ (square) feasible peg insertion solutions.
This first result confirms that—given a certain assembly strategy common across all pegs—multiple impedance solutions exist for the same peg. This finding suggests that there is a vast space of ”good-enough“ solutions for a given task, which contrasts with optimization methods that produce only a single set of impedance parameters [17, 49].
III-A2 Success Rate Distribution
For each of the pegs, the distribution of success rates for each impedance parameter ($k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}$) was computed by dividing each impedance parameter into a number of equally spaced bins and computing the success rate for each bin. Fig. 3 shows the observed distributions. A linear model was fit to each success rate distribution in each of the impedance parameters and for each peg. The coefficient of determination $R^{2}$ was computed, and the resulting linear fit was statistically tested against the degenerate null slope.
Interestingly, different pegs shared similar success rate distributions. The success rate grew with increase in vertical stiffness $k_{z}$ in all pegs, and the linear models presented good ($R^{2}>0.6$) coefficients of determination ($R^{2}_{triangle}=0.86,R^{2}_{square}=0.91,R^{2}_{hexagon}=0.77,R^{2}_{cylinder}=0.65$) with significant slopes ($p<0.05$) for all pegs. A similar behavior was observed for the success rate and the $k_{C}$ rotational stiffness across all pegs. The linear models presented high coefficients of determination ($R^{2}_{triangle}=0.73,R^{2}_{square}=0.85,R^{2}_{hexagon}=0.94,R^{2}_{cylinder}=0.88$) and significant slopes ($p<0.05$) for all pegs.
Another interesting finding was that the success rate appeared to be independent of the damping ratio parameters ($\zeta_{t},\zeta_{r}$). In fact, linear fits of the success rate to these two parameters presented low coefficients of determination and non-significant slopes ($p>0.05$) for all but two cases ($\zeta_{r_{square}},\zeta_{t_{hexagon}}$). Tab. II presents the complete details of the linear model results.
The success rate for the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$) was less consistent across pegs. Noticeably, the success rate distributions over $k_{x},k_{y}$ in the triangle peg and the cylinder peg were similar, with a concentration of success for low values of these two impedance parameters. However, the same was not true for the square and hexagon pegs. Appendix D contains the complete details of the linear model fitting performance on the other four impedance parameters ($k_{x},k_{y},k_{A},k_{B}$).
Figure 3: Distribution of success rates (blue bars) for each impedance parameter fpr the four different pegs (triangle, square, hexagon, and cylinder). For each histogram, the red solid lines represent the best linear fit, while the red dashed lines show the confidence intervals. An asterisk (*) was added for those linear fits that presented a significant slope.
TABLE II: Linear model fitting performance of the success rate distribution for four impedance parameters across all pegs. $R^{2}$ is the coefficient of determination of the linear fit. The p-value is the outcome of a statistical test of the linear model against the null slope. Asterisks denote significant differences.
Peg $k_{z}$ $k_{C}$ $\zeta_{t}$ $\zeta_{r}$
$R^{2}$ p-value $R^{2}$ p-value $R^{2}$ p-value $R^{2}$ p-value
Triangle 0.86 <0.001* 0.73 0.002* 0.10 0.501 0.52 0.066
Square 0.91 <0.001* 0.85 <0.001* 0.30 0.201 0.73 0.015*
Hexagon 0.77 <0.001* 0.94 <0.001* 0.58 0.048* 0.26 0.247
Cylinder 0.65 0.005* 0.88 <0.001* 0.14 0.403 0.09 0.524
III-B Claim 2: Lower dimensional representation of solutions
III-B1 Principal Component Analysis
To test whether there existed a non-arbitrary lower dimension subspace to represent the solutions (Claim 2), we projected the impedance solutions into a lower dimensional subspace using PCA. We found that 3 PCs could reasonably describe the data, a substantial reduction from the 8 dimensional solution space. Specifically, 3 PCs accounted for 83.8%, 96.6%, 95.9%, 88.7%, and 95.7% of the successful trials of the cylinder, hex, square, triangle, and all peg shapes, respectively. The amount of variance-accounted-for by each PC in the solution space is shown in fig. 4.
Figure 4: Top Row: Variance-accounted-for of the successful trials after Principal Component Analysis. Bottom Row: Variance-accounted-for of a random subset of trials after Principal Component Analysis. The black line represents a cumulative variance-accounted-for. The dashed red line represents a linear fit to the first three principal components; its slope is written in the legend.
To show that the reduction of the solution space was not a random artifact, we compared the slope of the variance-accounted-for by the first three principal components to the slope of the variance-accounted-for by a random subset of $n$ trials from the data, where $n$ is the number solutions reported in tab. I. This is shown in fig. 4. Then we conducted a pair-wise t-test between the variance-accounted-for slopes of the successful trials and the random trials. We found that the slopes of the successful trials were significantly steeper; that is, the slopes of the successful trials were significantly more negative ($p<0.001$). This indicates that the PCs extracted from the successful trials are not random artifacts; rather, they represent an impedance solution subspace.
Because the data patterns were effectively captured by three PCs, we were able to visualize them in a 3-dimensional space without losing significant information. By projecting the failure trials into this 3-dimensional subspace, we observed that the failures frequently fell outside the identified solution subspace, particularly in the cylinder and triangle trials (fig. 5). This visualization underscores the existence of a meaningful, lower-dimensional solution space, reinforcing Claim 2: there is a non-arbitrary lower-dimensional space representing the impedance solutions. Moreover, analyzing this solution space highlighted the multitude of successful impedance parameter configurations within EDA, illustrating the flexibility and adaptability of this approach. This will be further elaborated in sec. III-D.
III-C Claim 3: Different assembly strategies for individual pegs
III-C1 K-means Clustering
Figure 5: Left Column: The successful (green) and unsuccessful (red) trials of the cylinder, hex, square, and triangle pegs are projected into the solution space obtained using PCA. Note that often the failures fall outside the PCA-reduced solution space. Right Column: The coefficients of each of the three Principal Components are shown.
To test Claim 3—families of solutions exist which encode different assembly strategies—we used k-means clustering to identify clusters of successful trials in our eight-dimensional impedance solution space. Specifically, we used the silhouette coefficient to identify the “optimal” number of clusters. In fig. 6 we see that the optimal number of clusters was four, with a silhouette coefficient of 0.33. Further, in fig. 6 we see that often multiple clusters contained different pegs. Only Cluster 1 (red) contained all four pegs; it was the only cluster to include cylinder solutions. Together, these results suggest that similar strategies enabled different tasks (different pegs), which supports Claim 3.
To verify the distinctiveness of the identified clusters, we performed a one-way Multivariate Analysis of Variance (MANOVA) with cluster group as the independent variable, having four levels. The MANOVA results revealed significant differences among the clusters on the combined dependent variables, suggesting that the clusters’ multivariate means occupied at least three dimensions:
  • Dimension 0: Wilks’ Lambda $=0.0315$, $\chi^{2}(24)=3693.09$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.2383$, $\chi^{2}(14)=1,531.56$, $p<0.001$
  • Dimension 2: Wilks’ Lambda $=0.9824$, $\chi^{2}(6)=18.91$, $p=0.0043$
Overall, the MANOVA results validated the clusters identified by k-means clustering, indicating distinct multivariate mean vectors for each cluster. This supports Claim 3, that similar strategies facilitated different tasks (i.e., different pegs).
Figure 6: Top Left: The successful parameters of all pegs. Top right: the silhouette coefficient found from k-means clustering as a function of the number of clusters. An optimal number of clusters is denoted by a red X. Bottom: The successful trials of all pegs were projected into their solution space obtained using PCA. Each color denotes a different cluster identified by the k-means clustering algorithm. In the legend, “Cl” is an abbreviation for cluster. Different markers’ shape denotes the different pegs.
Figure 7: Left Column: The successful trials of each peg are projected into their solution space obtained using PCA (see fig. 5). Each color denotes a different cluster identified by the k-means clustering algorithm. Right Column: The successful parameters of each peg.
When using k-means clustering to identify families of solutions in the cylinder, hex, square, and triangle pegs we found 2, 4, 3, and 2 solution clusters, respectively. These clusters had silhouette coefficients of 0.33, 0.41, 0.29, and 0.37 for the cylinder, hex, square, and triangle pegs respectively (appendix E). The right column of fig. 7 informs the different solution strategies that were identified. For example, in the successful trials of the cylinder peg, Cluster 1 contained lower values of $k_{z}$ while Cluster 2 contained higher values of $k_{z}$. The difference between the 4 clusters in the successful hex trials depended on the combination of $k_{x}$ and $k_{y}$: Cluster 1 contained higher values of $k_{x}$ and lower values of $k_{y}$, Cluster 2 contained higher values of both $k_{x}$ and $k_{y}$, Cluster 3 contained lower values of $k_{x}$ and higher values of $k_{y}$, and Cluster 4 contained lower values of both $k_{x}$ and $k_{y}$. The three clusters identified in the successful square trials contained combinations of values in $k_{x}$, $k_{y}$, and $k_{z}$. Finally, the two clusters identified in the successful triangle trials differed in their range of values in $k_{y}$. All in all, this indicates that through the constrained assembly strategy multiple combinations of $k_{x}$ and $k_{y}$ were feasible. These results support Claim 3: families of solutions exist which encode different assembly strategies.
To determine if the identified clusters were indeed distinct, we conducted a one-way Multivariate Analysis of Variance (MANOVA) on the identified clusters. For the Cylinder, the independent variable was the cluster group with two levels. The MANOVA results showed significant differences among the clusters on the combined dependent variables, indicating that the multivariate means of the clusters lay in a space of 1 dimension (Wilks’ Lambda $=0.2198$, $\chi^{2}(8)=213.62$, $p<0.001$).
For the Hex, the independent variable was the cluster group with four levels. The MANOVA results showed significant differences among the clusters on the combined dependent variables. The results indicated that the multivariate means of the clusters lay in a space of at least three dimensions:
  • Dimension 0: Wilks’ Lambda $=0.0197$, $\chi^{2}(24)=683.51$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.1854$, $\chi^{2}(14)=293.27$, $p<0.001$
  • Dimension 2: Wilks’ Lambda $=0.9087$, $\chi^{2}(6)=16.65$, $p=0.0106$
For the Square, the independent variable was the cluster group with three levels. The MANOVA results showed significant differences among the clusters on the combined dependent variables. The results indicated that the multivariate means of the clusters lay in a space of at least two dimensions:
  • Dimension 0: Wilks’ Lambda $=0.1029$, $\chi^{2}(16)=1454.05.39$, $p<0.001$
  • Dimension 1: Wilks’ Lambda $=0.4322$, $\chi^{2}(7)=536.42$, $p<0.001$
For the Triangle, the independent variable was the cluster group with two levels. The MANOVA results showed significant differences among the clusters on the combined dependent variables, indicating that the multivariate means of the clusters lay in a space of 1 dimension (Wilks’ Lambda $=0.2391$, $\chi^{2}(8)=135.91$, $p<0.001$).
In summary, the MANOVA results consistently supported the validity of the clusters identified by k-means clustering across different shapes, indicating distinct multivariate mean vectors for each shape’s cluster groups. We conclude that families of ”good enough“ solutions exist which encode different assembly strategies.
III-D Claim 4: Common assembly strategies across all pegs
III-D1 Neural Networks for Binary Classification
As we have shown in our results, a range of successful parameters exist for each peg (Claim 1). These parameters might be grouped into different assembly strategies for each individual peg (Claim 3). To examine whether a common assembly strategy existed across all pegs (Claim 4), we created a binary classification neural network to sample over the whole space of successful parameters (sec. II).
[…]
We used the model above to produce 100,000 sets of successful impedance parameters; these data will be referred to as the predicted solutions. A PCA was conducted on a combined dataset of the 100,000 predicted solutions and the previously found successful trials. The projection into the first three PCs is shown in fig. 8 (note that the clusters here are the same as those in fig. 6). Three dominant PCs were present, accounting for 96.7% of the variance, while subsequent PCs accounted for less than 3% each. A few interesting findings can be observed. First, the predicted solution space for all pegs (shown in grey) is in fact not a uniform cube, suggesting that there is a subspace—or a family of solutions—common to all pegs. Note that areas in the solution cube outside of this subspace suggest impedance solutions that are not common to all pegs. For example, Cluster 1 (red) contains a solution strategy encompassing all four pegs. The predicted solution space here contains a higher density of predicted solutions in this area (grey) as opposed to other areas (e.g., Cluster 4); this is best seen in the top right of fig. 8. This intuition is consistent with our results.
Figure 9 presents the predicted success rates for all peg types, based on the same 100,000 data points. Although it is challenging to interpret all solutions due to the neural network predicting outcomes across all pegs, some interesting trends can be observed. The success rate was highest for low linear stiffness values $k_{x}$ and $k_{y}$, suggesting that the peg more easily slips into the hole during assembly with these lower parameters, compensating for imperfect alignment. A similar trend to that observed in fig. 3 is noted: success increases with higher vertical stiffness $k_{z}$ and rotational stiffness $k_{C}$. Once the peg is aligned with the hole, a certain force must be exerted, hence higher values of $k_{z}$ are more successful. Additionally, the original programmed orientation about the x-axis seemed to be similar across all pegs, which may explain the higher success rates for high rotational stiffness $k_{C}$. Interestingly, $k_{A}$ appears to be successful across all parameter ranges. The rotational stiffness $k_{B}$ shows the highest success rates at medium values, which might indicate that the final orientation of the peg and hole was not perfectly programmed and varied slightly across pegs. This could explain the optimal balance of maintaining orientation without being overly stiff about the y-axis. The damping parameters $\zeta_{t}$ and $\zeta_{r}$ were successful across all ranges, with a growing success rate for increasing values of $\zeta_{t}$ and a declining success rate for decreasing values of $\zeta_{r}$. Still, we conclude that for our assembly application, the effect of the selected damping ratio was minor, most likely due to the low speed of the task.
Figure 8: Left: The successful parameters of all pegs (from Fig. 6) as well as the predicted solutions projected into a common PC space. Right: Various 2D views of the 3D image.
Figure 9: Distribution of success rates for parameters predicted as successful by the neural network. The distributions were generated by sampling impedance parameters uniformly between the minimum and maximum values specified in sec. II until 100,000 successes were achieved, resulting in a success rate of $4.32\%$.
III-E Application: Success prediction for peg-in-hole assembly
Even for a given assembly strategy (submovements and oscillations in algorithm 1), determining a set of feasible impedance parameters $\{k_{x},k_{y},k_{z},k_{C},k_{B},k_{A},\zeta_{t},\zeta_{r}\}$ is not trivial [50]. Hence, it requires a lot of training and trial-and-error for robot programmers. For users with pegs similar to this work, the trained neural network (sec. III-D) can be used as a success predictor. We provide Python code on Github [^5] https://github.com/mosesnah-shared/ImpedanceClassification/blob/main/ImpedanceClassification/success_failure_predictor.py for the user to input a set of impedance parameters; the neural net can then predict the result. This is especially beneficial for inexperienced robot programmers, as they get immediate feedback without trial-and-error procedures on a real robot.
To evaluate the outcome of the neural net, we randomly selected 30 predicted solutions and tested these parameter sets for each peg on the real robot. All parameter sets were successful except for one parameter set each for the square and triangle pegs, respectively. Namely, these solutions were successful in 118 out of 120 trials. A video of the experiments on the robot can be found on Dropbox [^6] https://www.dropbox.com/scl/fo/gav5l7kpzqs69dzbi06xj/AMGmfsqsbLDaThPr7xVikWo?rlkey=pc443f61levpdeme9exejhsa1&dl=0.
IV Discussion
Our study has successfully confirmed the claims we set out to investigate, providing valuable insights into the use of Elementary Dynamic Actions (EDA) in robotic peg-in-hole assembly tasks:
  • Claim 1: Success with a range of parameters
  • Claim 2: Lower dimensional representation of solutions
  • Claim 3: Different assembly strategies for individual pegs
  • Claim 4: Common assembly strategies across all pegs.
Claim 1: We have demonstrated that EDA enables a range of successful parameters, rather than being limited to a single optimal set. This flexibility contrasts with traditional optimization methods, which often focus on finding one optimal solution. Our findings show that multiple ”good enough“ impedance settings can lead to successful assembly, highlighting the adaptability of EDA for managing physical interactions.
Claim 2: Our analysis revealed that there exists a non-arbitrary lower-dimensional space to represent successful solutions, as evidenced by the principal components (PCs) derived from our data. This dimensionality reduction simplifies the complexity of the parameter space, making it easier to identify and utilize effective impedance settings for assembly tasks.
Claim 3: We identified families of solutions that encode different assembly strategies, supporting the notion that similar strategies can enable different tasks, such as the assembly of various peg shapes. This finding underscores the versatility of EDA in accommodating task-specific variations while maintaining overall effectiveness.
Claim 4: Our findings demonstrate that certain impedance settings can be generalized across different peg types, indicating the presence of common assembly strategies. This generalization is particularly significant as it suggests that despite the variability in peg shapes and sizes, there are underlying principles of impedance parameters that can be applied. The ability to identify and utilize these common strategies can streamline programming efforts, reduce the need for extensive parameter tuning, and enhance the overall effort of creating robotic assembly processes.
The overall findings from these claims demonstrate the robustness and flexibility of EDA in handling complex robotic assembly tasks. The ability to use a range of successful impedance parameters (Claim 1) shows that EDA allows for multiple solutions rather than being confined to a single optimal set, which is crucial for adaptability in dynamic environments. The identification of a lower-dimensional space for these solutions (Claim 2) simplifies the complexity of the parameter space, making it easier to identify effective settings and facilitating more efficient programming and tuning processes. Furthermore, the discovery of different assembly strategies for individual pegs (Claim 3) highlights EDA’s capacity to adapt to specific task requirements, ensuring that robots can handle a variety of shapes and assembly challenges with tailored approaches. Additionally, the presence of common assembly strategies across all pegs (Claim 4) underscores the potential for generalization, enabling the application of universal principles of impedance parameters across different tasks. This not only streamlines the programming process but also reduces the need for extensive parameter tuning, enhancing the overall efficiency of robotic assembly operations.
By generalizing impedance settings into categories like low, medium, and high, the programming process becomes more intuitive and accessible. This modularity and categorization support efficient programming and quick adaptation to small offsets in peg locations, aligning with Bernstein’s [31] concept of “repetition without repetition,” where humans adapt their movements to environmental variations.
As all of the CAD data and software of this work are publicly available, we hope to enable researchers to test the data set, extend it, or even use it for real peg-in-hole assembly applications in industry.
IV-A Takeaway messages for robotics
This study presents several key insights into the control of robotic peg-in-hole assembly tasks, particularly through the modulation of impedance parameters. Here are the main takeaways of our work for roboticists working in assembly and control:
  • Reconsidering the need for “high control gains” in contact-rich manipulation: While conventional strategies in robotics control often prioritize higher stiffness values to improve precision and stability, our findings suggest that lower stiffness values, especially in terms of $k_{x}$ (translation along $x$), $k_{y}$ (translation along $y$), $k_{A}$ (rotation about $z$), $\zeta_{t}$ (translational damping ratio), and $\zeta_{r}$ (rotational damping ratio), are effective for successful peg-in-hole assembly tasks. This makes sense from a physical perspective, as the peg can slip into the hole more easy for low translational impedance parameters. This challenges the current trend of focusing on faster control loops to increase control values and suggests a potential shift towards more compliant interaction strategies in contact-rich tasks.
  • Generalizing impedance settings for programming interfaces: The ability to generalize impedance regions for all assembly tasks could greatly simplify robotic programming interfaces. By categorizing impedance settings into simple groups like low, medium, or high, programmers can more intuitively select appropriate settings, enhancing the accessibility and efficiency of robot programming for assembly tasks.
  • Rethinking the role of damping ratio for contact-rich tasks: Our study indicates that the influence of damping ratio on the success of assembly tasks might be less critical than previously assumed. Given this finding, it may be practical to adopt a default value for damping ratio parameters in most scenarios, simplifying the tuning process without significantly compromising performance. Of course, more dynamic tasks (i.e., with higher speeds) with less physical constraints might be more dependent on the damping ratio.
  • Optimal solutions and unconstrained directions: Consistent with other studies focused on contact-rich tasks, our research confirms that increasing impedance parameters along unconstrained directions, such as $k_{z}$ in our case, increases successful outcomes [23, 51]. However, for all other parameters, smaller to medium impedance parameters were found to be equally effective, suggesting that the optimal set of impedance values may not always trend towards the higher end of the spectrum. This insight encourages a more balanced approach to setting impedance in robotic control, where the specific task characteristics guide the parameter choices rather than a one-size-fits-all strategy.
By reconsidering traditional approaches to choose impedance values in robotic control, these insights aim to enhance the flexibility and effectiveness of robots performing complex assembly tasks. This could lead to more adaptable and user-friendly robotic systems that are capable of handling a variety of assembly operations with greater efficiency and less extensive tuning.
V Limitations
Our study also indicates that the effectiveness of EDA is not only dependent on the impedance parameter choice but also dependent on the chosen strategy for kinematic primitives. While EDA’s modularity offers many benefits, the specific combination of submovements and oscillations used in our experiments played a crucial role in achieving successful outcomes. This dependency suggests that further refinement and customization of kinematic strategies may be necessary to optimize EDA’s performance across different tasks. Here, teaching by demonstration might be a promising way to facilitate the programming of kinematic primitives [27].
Our study involved several parameter choices that could influence the outcomes of our investigation. For instance, during the Principal Component Analysis, we reduced the dimensionality of the solution space from eight to three principal components, capturing a certain percentage of the variance. This choice, while effective for our purposes, may not generalize across different contexts or datasets.
In our analysis, we employed k-means clustering in conjunction with silhouette coefficients to ascertain the optimal number of clusters. However, it is important to acknowledge that the silhouette coefficient is fraught with several limitations. These limitations include its sensitivity to the shape and density of clusters, as it assumes clusters to be spherical and evenly sized, which may not always be the case in real-world data. In our case, the resulting silhouette coefficients were relatively low, ranging from 0.22 to 0.41, suggesting a possible absence of a distinct separation between clusters. This highlights one of the inherent challenges in clustering analysis: the difficulty of defining a universally accepted threshold for silhouette coefficients that reliably demarcates well-defined clusters. Determining the appropriate silhouette coefficient for effective cluster delineation remains an active and evolving area of research, which is beyond the scope of this paper.
In the neural network model for binary classification, we used a structure with two hidden layers, each consisting of 64 neurons, and trained the network over $2^{13}$ epochs. The Adam optimizer with a fixed learning rate of $1\times 10^{-4}$ was used to update the weights. Although these are common choices in neural network design, other configurations and hyper parameters could also have been effective in producing a feasible neural network.
An additional limitation of our study is the lack of coupling between degrees of freedom. Traditional solutions for peg-in-hole insertion, such as remote-center compliance (RCC), inherently couple the stiffness of different degrees of freedom to facilitate alignment and insertion tasks [52, 53]. Our simplification to independent impedance parameters was made to keep the system accessible for novice programmers, aiming to maintain simplicity in implementation. However, it is important to acknowledge that this decoupling may overlook some of the benefits provided by more sophisticated compliance mechanisms. Future work should investigate the integration of coupled impedance parameters to enhance the robustness and accuracy of EDA in more complex assembly tasks.
Finally, while our results are promising for peg-in-hole assembly, further investigations are needed to evaluate how our results may be generalizable to other manipulation tasks. Different tasks may present unique challenges that require further investigation to determine the applicability and adaptability of EDA in those contexts. Future research should explore the use of EDA in a broader range of assembly operations to validate its generalizability and effectiveness in various industrial applications.
Appendix A Motor primitives of EDA
Consider a system with $n$ degrees of freedom (DOF).
Figure 10: (A) The three primitives of Elementary Dynamic Actions (EDA). Submovements and oscillations are kinematic primitives used to generate discrete and periodic movements respectively [35], and mechanical impedance is an interactive primitive used to handle physical interaction [36, 54, 55, 56, 57]. (B) The three primitives are combined using a Norton equivalent network model.
A-1 Submovement
A submovement $\mathbf{x}_{0}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ is a smooth trajectory in which its time derivative is a unimodal function, i.e., has a single peak value:
$\dot{\mathbf{x}}_{0}(t)=\mathbf{v}\ \hat{\sigma}(t)$ (6)
In this equation, $\hat{\sigma}:\mathbb{R}_{\geq 0}\rightarrow[0,1]$ denotes a smooth unimodal basis function with peak value 1; $\mathbf{v}\in\mathbb{R}^{n}$ is the velocity amplitude of the submovement. Submovements model discrete reaching motions, and therefore $\hat{\sigma}(t)$ has a finite support, i.e., there exists $T>0$ such that $\hat{\sigma}(t)=0$ for $t\geq T$. [58]
A-2 Oscillation
An oscillation $\mathbf{x}_{0}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}^{n}$ is a smooth non-zero trajectory which is a periodic function:
$\forall t>0:~{}~{}\exists T>0:~{}~{}\mathbf{x}_{0}(t)=\mathbf{x}_{0}(t+T)$ (7)
Compared to submovements, oscillations model rhythmic and repetitive motions.
A-3 Mechanical Impedance
A mechanical impedance $\mathbf{Z}:\Delta\mathbf{x}(t)\longrightarrow\mathbf{F}(t)$ is an operator which maps (generalized) displacement $\Delta\mathbf{x}(t)\in\mathbb{R}^{n}$ to (generalized) force $\mathbf{F}(t)\in\mathbb{R}^{n}$ [36]. Here, $\Delta\mathbf{x}(t)\equiv\mathbf{x}_{0}(t)-\mathbf{x}(t)$ is the displacement between an actual trajectory of (generalized) position $\mathbf{x}(t)$ and a virtual trajectory $\mathbf{x}_{0}(t)$ to which the mechanical impedance is connected.
The dynamics of physical interaction can be controlled by modulating mechanical impedance. For instance, tactile exploration and manipulation of fragile objects should evoke the use of low stiffness, while tasks such as drilling a hole require high stiffness for object stabilization [54]. Under the assumption that the environment is an admittance, mechanical impedance can be linearly superimposed even though each mechanical impedance is a nonlinear operator [50, 36]: $\mathbf{Z}=\sum\mathbf{Z}_{i}$.
A-4 Norton Equivalent Network Model
The three distinct motor primitives of EDA (fig. 10A) can be combined using a Norton equivalent network to model physical interaction [36] (fig. 10B). The forward-path dynamics specify the virtual trajectory $\mathbf{x}_{0}(t)$, which consists of submovements and/or oscillations. The interactive dynamics, which consists of mechanical impedance $\mathbf{Z}$, determines the generalized force output $\mathbf{F}(t)$. The generalized displacement input is denoted $\Delta\mathbf{x}(t)$. Hence, a key objective of EDA is to find appropriate choices of $\mathbf{x}_{0}(t)$ and $\mathbf{Z}$ to produce the desired interactive behavior at the interaction port $(\mathbf{F},\dot{\mathbf{x}})$ (fig. 10B). This approach stands as an alternative to traditional robotic control, which predominantly focuses on motion control. Instead, it emphasizes the importance of interactive dynamics in achieving desired behavior. This paper will choose a fixed strategy for the zero-force trajectory, $\mathbf{x}_{0}(t)$, and investigate the influence of different choices of impedance parameters, $\mathbf{Z}$.
Appendix B Categories and parameter ranges
TABLE III: Three different categories for each of the parameter ranges. The units for the translational stiffnesses and rotational stiffnesses are N/m and Nm/rad, respectively. The damping ratio is unitless.
Category $k_{t}$ $k_{r}$ $\zeta$
Small $50\leq k_{t}\leq 300$ $5\leq k_{r}\leq 10$ $0.1\leq\zeta\leq 0.3$
Medium $301\leq k_{t}\leq 700$ $11\leq k_{r}\leq 80$ $0.4\leq\zeta\leq 0.7$
Large $701\leq k_{t}\leq 1000$ $81\leq k_{r}\leq 200$ $0.8\leq\zeta\leq 0.9$
Appendix C MANOVA to identify distinct clusters
To determine if the identified clusters were indeed distinct, we conducted a one-way Multivariate Analysis of Variance (MANOVA). MANOVA tests the null hypothesis that the means of each group are the same $n$-dimensional multivariate vector and that any observed differences in the sample are due to random chance. Note that the null hypothesis is examined across each of the $n$ dimensions where the largest possible dimension is either the dimension of the space or one less than the number of groups. Specifically, we used MANOVA to assess the effect of cluster group on the eight impedance parameters (dependent variables) for different shapes (cylinder, hex, square, and triangle). The independent variable was the cluster group, which had varying levels depending on the number of clusters identified for a given shape. The MANOVA was implemented using the manova1 function in MATLAB [59]. This function returns an estimate of the dimension of the space containing the group means and tests the null hypothesis that the group means lie within a subspace of the given dimension. Specifically, it assesses whether the means of the groups can be distinguished when projected into subspaces of increasing dimensionality, ranging from dimension zero (indicating no significant difference between the group means) up to dimension $k-1$, where $k$ is the number of groups. The upper limit of $k-1$ dimensions arises because there are $k-1$ degrees of freedom in comparing $k$ groups. The results from the manova1 function were used to determine the significance of the differences among the clusters.
Appendix D Linear regression models
Tab. IV reports the coefficients of determination and related p-values for the linear models used to fit the relative success rates of fig. 3.
TABLE IV: Coefficients of determination ($R^{2}$) and p-values of the linear fit of each relative success histogram for the four different pegs. Asterisks (*) denote significant difference between the linear models and the degenerate null slope.
Triangle Square Hexagon Cylinder
$R^{2}$ p-value $R^{2}$ p-value $R^{2}$ p-value $R^{2}$ p-value
$k_{x}$ 0.64 0.005* 0.60 0.009* 0.08 0.432 0.64 0.006*
$k_{y}$ 0.77 <0.001<0.001*< 0.001 ∗ 0.32 0.088* 0.02 0.663 0.65 0.005*
$k_{z}$ 0.86 <0.001<0.001*< 0.001 ∗ 0.91 <0.001<0.001*< 0.001 ∗ 0.77 <0.001<0.001*< 0.001 ∗ 0.65 0.005*
$k_{A}$(z) 0.54 0.016* 0.22 0.174 0.01 0.804 0.52 0.019*
$k_{B}$(y) 0.42 0.043* 0.48 0.026* 0.14 0.281 0.05 0.542
$k_{C}$(x) 0.73 0.002* 0.85 <0.001<0.001*< 0.001 ∗ 0.94 <0.001<0.001*< 0.001 ∗ 0.88 <0.001<0.001*< 0.001 ∗
$\zeta_{t}$ 0.10 0.501 0.30 0.201 0.58 0.048* 0.14 0.403
$\zeta_{r}$ 0.52 0.066 0.73 0.015* 0.26 0.247 0.09 0.524
Figure 11 shows the distribution of success rates for 4 different neural networks specifically trained on each individual peg: triangle, square, hexagon, and cylinder. The individual neural networks shared the same structure of the neural network presented in the main manuscript. They were trained using the same split (70/30) between training and validation data. The resulting accuracy of the neural networks were respectively: triangle $98.53\%$, square $95.84\%$, hexagon $97.41\%$, and cylinder $98.83\%$. The success rate distributions were generated by uniformly sampling the 8 impedance parameters between their maximum and minimum values until 100,000 successes were collected. The resulting success rate for each peg was: square $s_{r}=27.07\%$, hexagon $s_{r}=5.87\%$, triangle $s_{r}=1.02\%$, cylinder $s_{r}=2.36\%$. It is interesting to observe how in some cases (square, hexagon) the success rate of the neural network was higher than the experimental data, while in the other cases (triangle, cylinder) the success rate was practically indistinguishable from the experimental data.
Figure 11: Distribution of success rates for the parameters that were predicted to be successful according to the trained neural network. The distributions were generated by uniformly sampling impedance parameters between the prescribed minimum and maximum values presented in Section II until 100,000 successes were generated. The resulting success rate (number of successful parameters over number of tested parameters) for each peg were: triangle $s_{r}=1.02\%$, square $s_{r}=27.07\%$, hexagon $s_{r}=5.87\%$, cylinder $s_{r}=2.36\%$.
Appendix E Silhouette coefficients
Figure 12: The silhouette coefficient found from k-means clustering as a function of the number of clusters. The optimal number of clusters for each peg is denoted by the red X.
Acknowledgment
This work was supported in part by the MIT/SUSTech Centers for Mechanical Engineering Research and Education. MCN was supported in part by a Mathworks Fellowship. JL was supported in part by the MIT-Novo Nordisk Artificial Intelligence Postdoctoral Fellows Program. AMWJ was supported by the Accenture Fellowship.
We want to thank KUKA Deutschland GmbH for providing the LBR iiwa robots.